OpenStudy (anonymous):

What is the value of x? A right angle is shown divided into two parts. The measure of one part of the right angle is 30 degrees. The measure of the other part is 4x 15 30 45 60

Mehek (mehek14):

do you have a pic of it?

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

The answer is 4x = 90-30 4x/ 4 = 60 / 4 x = 15 The answer is 15